### v2

# SynthVLM: High-Efficiency and High-Quality Synthetic Data for Vision Language Models

Abstract.
Recently, with the rise of web images, managing and understanding large-scale image datasets has become increasingly important. Vision Large Language Models (VLLMs) have recently emerged due to their robust vision-understanding capabilities. However, training these models requires vast amounts of data, posing challenges to efficiency, effectiveness, data quality, and privacy. In this paper, we introduce SynthVLM, a novel data synthesis pipeline for VLLMs. Unlike existing methods that generate captions from images, SynthVLM employs advanced diffusion models and high-quality captions to automatically generate and select high-resolution images from captions, creating precisely aligned image-text pairs. Leveraging these pairs, we achieve state-of-the-art (SoTA) performance on various vision question answering tasks, maintaining high alignment quality and preserving advanced language abilities. Moreover, SynthVLM surpasses traditional GPT-4 Vision-based caption generation methods in performance while significantly reducing computational overhead. Crucially, our method’s reliance on purely generated data ensures the preservation of privacy, achieving SoTA performance with just 100K data points (only 18% of the official dataset size). The code is made available at https://github.com/starriver030515/SynthVLM.
$\dagger$ The first two authors have equal contributions.
* Corresponding Author
1. Introduction
In recent years, with the rapid advancements in large language models (LLMs) (OpenAI, 2023a; Touvron et al., 2023) and multimodal large language models (MLLMs) (Zhao et al., 2023; Wu et al., 2023), data management has become a crucial aspect of these technologies (Fernandez et al., 2023; Trummer, 2023; Chen et al., 2023a; Miao et al., 2024; Nie et al., 2023). At the same time, Bai et al. (2024) also demonstrates that data processing, selection, and management can significantly influence the performance of MLLMs.
Among MLLMs, VLLMs achieve competitive performance in traditional multimodal tasks such as image classification (Chen et al., 2024), image understanding (Li et al., 2023b, c), and image captioning (Bai et al., 2023). Moreover, their excellent language understanding capabilities enable strong performance in text-rich tasks, such as vision question-answering (Liu et al., 2023b, a) and image-text retrieval (Chen et al., 2024).
While most existing VLLMs focus on modifying model architecture to utilize information from multiple modalities (Li et al., 2023b, c; Bai et al., 2023; Liu et al., 2023b, a; Chen et al., 2024), data also significantly impacts the success of VLLMs. For instance, Li et al. (2023d); Wang et al. (2024a) demonstrates that higher-quality training data can enhance the performance of VLLMs. The key to ensuring high-quality data lies in the precise alignment between multimodal data, such as the alignment between captions and images. With the introduction of the DataComp (Gadre et al., 2023), increasing research efforts have focused on exploring how to achieve effective caption-image alignment. Key approaches include the use of advanced alignment techniques, the development of refined annotation guidelines, and the implementation of novel training methodologies.
With the advancement of generative models, data generation strategies have increasingly been utilized to achieve data creation and alignment. For example, Nguyen et al. (2023) employed BLIP2 to generate numerous image captions, achieving SoTA results on DataComp. In the domain of VLLMs, Chen et al. (2023b) utilized GPT-4 Vision to produce highly descriptive image captions, leading to significant improvements in LLaVA. The integration of these generative models has opened new avenues for enhancing data quality and alignment, further boosting VLLM performance.
Despite the notable contributions and advances in VLLMs, current data generation and alignment strategies for VLLMs ignore generating images and face the following three key challenges:
Figure 1. We compared the synthetic image-text dataset with existing datasets. As shown in (a), the generated image can avoid content such as watermarks and advertisements. In (b), the caption mentions a joicy named plate, however in the bottom right image, joicy is actually a book. The generated images better reflect the content of the captions. Additionally, the resolution of the generated images is 1024x1024, which is higher than the existing images, is more beneficial for model training and expansion.
C1. Low Data Quality. Our evaluations using metrics such as CLIPScore  (Hessel et al., 2021) and training results on VLLMs reveal that existing datasets still align the modalities sub-optimally. Web images often suffer from issues such as blurriness and watermarks. Additionally, methods employing BLIP2  (Nguyen et al., 2023) for generating captions frequently result in logical inconsistencies and unclear grammar, which mislead VLLM training and degrade their language capabilities.
C2.Poor Effectiveness. It has been demonstrated in (Du et al., 2023; He et al., 2020; Chen et al., 2023c) that low-quality data can lead to poor model performance. Consequently, VLLMs are usually trained on low-quality data, the VLLMs exhibit reduced effectiveness.
C3. Low Efficiency. Methods that rely on manual captioning are labor-intensive and resource-demanding. Automated solutions like ShareGPT4v  (Chen et al., 2023b), which employ GPT-4 Vision for labeling, are expensive and difficult to scale. Moreover, current strategies often necessitate the creation of extensive datasets to enhance performance, leading to significant data redundancy.
C4. Security Risks. Utilizing internet-sourced data introduces numerous security and privacy concerns  (Li et al., 2023a; Das et al., 2024). They may contain personal information or copyrighted materials, posing potential legal and ethical challenges. Moreover, the inclusion of sensitive or inappropriate content within training datasets can instigate ethical issues, thereby compromising the models’ integrity and fairness.
To address these issues, we introduced a new data generation pipeline: caption-to-image synthesis. We first implemented a quality selection process for high-quality caption data. Subsequently, we employed advanced diffusion models to generate images from these captions. For quality control, we used CLIPScore as the quality metric to select high-quality image-text pairs. Our data generation method achieved higher alignment between images and captions compared to existing approaches. Utilizing 100K curated synthetic data, we achieved SoTA results on multiple benchmarks, using only 18% of the official dataset size.
Overall, our contributions are as follows:
Figure 2. With only 100k synthetic pre-training data, our SynthVLM outperforms the LLAVA 1.5 model, which is trained on 558k data.
  • New Perspective. To the best of our knowledge, we are the first to utilize Caption-to-Image strategy to construct a high-aligned training dataset for VLLMs and achieved superior performance with just 100K data points(only 18% of the official dataset size). This sets a precedent for utilizing generated images to train large-scale VLLMs, offering a solution to the limitation of data availability, quality, and privacy.
  • New Method. We proposed a new image-text pair generation pipeline to ensure high-quality data. Additionally, we proposed a new paradigm for utilizing generated images for VLLMs training.
  • SoTA Performance. (1) High Data Quality. As shown in Figure 1, our generated data achieved higher CLIPScores, indicating better image-text alignment. Additionally, we achieved higher resolution, which is beneficial for tasks requiring high-quality images. Furthermore, our generated images avoid issues such as blurriness and watermarks.
    (2) High Efficiency and Effectiveness. As shown in Figure 2, with only 18% of the high-quality pre-training data, our method outperforms the baseline that uses 100% of the data. We not only achieve SoTA performance on vision understanding tasks but also demonstrate excellent pure text task capabilities, highlighting superior modality alignment.
    (3) Data Privacy. Using generated data avoids the need for real personal images, documents, and other sensitive information, ensuring data privacy.
2. Related Work
2.1. Diffusion Model
Denoising diffusion probabilistic models (DDPMs)  (Ho et al., 2020; Rombach et al., 2022; Podell et al., 2023) are a class of generative models renowned for their ability to generate extremely high-quality images. The core idea of DDPMs involves modeling the data distribution by gradually adding Gaussian noise to the input image during the forward process and then predicting and removing this noise to reconstruct the image during the backward process.
Given a source image data distribution $x_{0}\sim q(x_{0})$, Gaussian noise is added over $T$ steps to obtain $x_{T}$. The forward process is defined as:
$q(x_{1},\ldots,x_{T}\mid x_{0}):=\prod_{t=1}^{T}q(x_{t}\mid x_{t-1}),$
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
where $\beta_{t}$ controls the variance of the noise added at each step.
The distribution after $t$ steps can be written as:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$.
The backward process aims to reconstruct the data by learning a series of Gaussian distributions that approximate the forward process:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are neural networks parameterized by $\theta$.
While DDPMs have shown promising results, several improvements have been proposed to enhance their efficiency  (Song et al., 2021; Zhu et al., 2024) and sample quality  (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021). The superior performance of diffusion models has been leveraged in various sub-tasks, including image generation, image translation, inpainting  (Lu et al., 2022; Su et al., 2023). Our approach leverages diffusion models to create high-quality caption-image pairs specifically for VLLM training. These generated pairs not only enrich the training data but also provide a more diverse and comprehensive dataset, enabling VLLMs to learn better representations and achieve higher performance in downstream tasks such as image captioning and visual question answering.
2.2. Vision Language Models
The integration of visual knowledge into large language models (LLMs) has become a pivotal area of research due to the rapid advancements in LLMs. VLLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks  (Liu et al., 2023d; Zhang et al., 2022; Li et al., 2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP  (Radford et al., 2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters  (Liu et al., 2023b, a; Li et al., 2023b, 2022a; Jian et al., 2023; Lu et al., 2023a) are introduced to further integrate different modalities. For example, LLaVA  (Liu et al., 2023b, a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP  (Li et al., 2022a, 2023b) utilize cross-attention to enhance modality integration.
Recent studies  (Wang et al., 2024b; Chen et al., 2023b; Liu et al., 2023b, a; Li et al., 2023f) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al., 2023b, a) and ShareGPT4V  (Chen et al., 2023b) have shown remarkable advancements in understanding and following complex instructions through instruction tuning. Although these improvements help align the vision modality and establish a solid basis for cross-modal comprehension, they require extensive datasets for training and could potentially diminish the model’s language capabilities. In this work, we propose a novel data generation strategy that leverages high-quality caption-image pairs, enabling the model to achieve state-of-the-art (SoTA) results with minimal data. By aligning during the pre-training (PT) phase, we have further enhanced the model’s language capabilities.
2.3. Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data. (Touvron et al., 2023; OpenAI, 2023b) In this scenario, the quality and quantity of data become paramount. LLMs, trained on vast amounts of data, can capture subtle nuances and complex patterns in language, excelling in various natural language processing tasks. However, the increase in data volume also brings new challenges, particularly in data management, cleaning, and annotation. (Bai et al., 2024) In this section, we mainly discuss the effectiveness of data quality and data selection.
Data Quality
: High-quality data can significantly enhance the performance of models (meta llama, 2024). As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation. (Bai et al., 2024) Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
Data Selection
: LLMs-based methods were commonly used in data selection. (Bai et al., 2024) For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023c) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023c) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023b) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
Given the critical role of data quality and selection in enhancing model performance, our paper focuses on leveraging advanced data selection techniques to optimize caption and image-text pair quality. By employing methods that integrate LLM data selection and image-text alignment scores, we aim to efficiently identify and utilize high-quality data for VLLMs.
2.4. Data Generation
Data has always been the key driver behind the success of LLMs. Recent advancements of LLMs largely due to the availability of large-scale, diverse, and high-quality datasets for training these models  (Lin et al., 2023). However, the scarcity of data and the high costs present substantial challenges in obtaining such datasets  (Xie et al., 2024; Xu et al., 2023a). Recent advancements in generating synthetic data and improving the performance of LLMs have shown promising results across various domains. Synthetic data holds great potential in building large-scale, high-quality datasets. Researchers have explored multiple approaches, from leveraging differential privacy to creating instruction-tuning frameworks, to enhance the quality, diversity, and utility of synthetic data  (Wang et al., 2024c; Wei et al., 2023; Lou et al., 2023; Yang et al., 2023). A key component in generating high-quality synthetic datasets is precise alignment. Fan et al.  (Fan et al., 2024) introduce REALIGN, a method that enhances the quality of instruction data by reformatting responses to better align with pre-established criteria and evidence, thereby improving LLMs’ alignment with human values while minimizing human annotation and model hallucinations. Li et al.  (Li et al., 2023e) build a high-quality instruction-following language model by automatically labeling human-written text with corresponding instructions and demonstrating highly effective self-alignment.
In the field of VLLMs, the task of constructing generative datasets has been relatively underexplored. VLLMs primarily focus on the alignment between images and captions. Existing approaches, such as ShareGPT4V  (Chen et al., 2023b), leverage GPT4-Vision to generate high-quality captions for images, thereby achieving alignment and producing SoTA results on models like LLaVA. However, this method incurs high costs and often results in suboptimal alignment due to the complexity of the captions. Our approach introduces a new method for aligning images and captions by utilizing text-to-image models to generate large-scale, high-quality data. Specifically, our method surpasses existing techniques in alignment accuracy and efficiency.
Figure 3. Curated image-text pair generation pipeline. We first use the diffusion model to generate images and then select the high clip score image-text pairs.
3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1 million caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. Then we meticulously select 100k high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM dataset with other existing datasets. In subsection 3.4, we summarize the VLLMs training pipeline by utilizing the synthetic dataset.
Figure 4. This is our process and prompt design for match assessment using GPT4-Vision. We consider various aspects, including the quality of the image and the match between the image and the caption, to help GPT4-Vision make a better selection. Based on this process, we compare SynthVLM with existing datasets from the model’s perspective.
3.1. Synthetic Dataset Construction
In this section, we introduce the image generation pipeline. First, we construct a large pool of captions for selection. Next, we generate high-quality images corresponding to these captions. We then select the best captions from the pool for image-text generation. Utilizing these high-quality captions, we employ diffusion models to generate the images.
Table 1. LCS abbreviates the LAION, CC, and SBU datasets. SynthVLM uses captions to generate images, while others use images to generate captions or manual labeling.
Name Image Source Caption Source Sample
COCO-Caption COCO Human 118K
BLIP-LCS LCS BLIP 558K
ShareGPT4V LCS, COCO, etc GPT4-Vision 100K
ShareGPT4V-PT LCS, COCO, etc Share-Captioner 1246K
SynthVLM Diffusion LCS, COCO, BLIP2-DataComp, etc 1000K
Data Source. To ensure the diversity of the captions, we combined human-generated and machine-generated captions. As shown in Table  1. The human-generated captions were primarily sourced from LAION, CC, and SBU, while the machine-generated captions were predominantly created using the method described in  (Nguyen et al., 2023). They utilize BLIP2 to regenerate captions for images in the datacomp dataset  (Gadre et al., 2023).
Caption Curation. To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. This filtering process was performed using GPT-3, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions.
The formula for calculating CLIPScore is as follows:
$\text{CLIPScore}(I,C)=\frac{\text{CLIP}(I)\cdot\text{CLIP}(C)}{||\text{CLIP}(I)||\cdot||\text{CLIP}(C)||}$
where $I$ represents the image, $C$ represents the caption, and $\text{CLIP}(I)$ and $\text{CLIP}(C)$ denote the image and text feature vectors extracted by the CLIP model. The dot product of the vectors is denoted by $\cdot$, and |||||$\cdot$||| | ⋅ | | denotes the norm of the vectors.
We selected the top 40% of image-caption pairs with the highest scores. These selected captions were included in the candidate caption set. At this step, we believe that captions with better image-caption alignment are more likely to generate high-quality, well-aligned images. Ultimately, we sampled a dataset of 1,000k captions for data generation. By using only captions, our method significantly reduces storage overhead and processing time. The caption curation pipeline is summarized in Figure 3(a).
Image Generation. After filtering through 1000k high-quality captions, we utilized Stable Diffusion XL (SDXL)  (Podell et al., 2023), a SoTA model which can efficiently generate high-quality, high-resolution images. By configuring SDXL’s steps to 60 and utilizing 8 A100 GPUs, we were able to generate all images within a week. Furthermore, SynthVLM generates images at a resolution of 1024x1024, effectively addressing the issue of low resolution found in existing datasets. This enhancement significantly improves the quality and utility of the training data for a variety of image generation and recognition tasks.
3.2. Synthetic Data Selection
In this section, we introduce the quality control process for synthetic datasets. To further ensure the alignment between images and their corresponding text descriptions, we employ CLIPScore a second time, evaluating the quality of image-text pairs with enhanced precision.
As shown in Figure 3(b), we initially computed CLIPScores for the 1, [^0] K synthetic image-caption pairs. We then selected the top 100K pairs that demonstrated the highest scores, indicating the most accurate and meaningful matches between images and captions. By curating this subset, we constructed a high-quality, highly aligned synthetic dataset.
Table 2. We compared the average CLIPScores of our synthetic dataset, ShareGPT4V, COCO-Caption, and BLIP-LCS. The results indicate that SynthVLM exhibits the highest alignment in terms of CLIPScore values.
Name Sample Avg CLIPScore
COCO-Caption 118K 0.31
BLIP-LCS 558K 0.32
ShareGPT4V 100K 0.32
SynthVLM 1000K 0.34
Curated-SynthVLM 100K 0.38
3.3. High Quality Synthetic Dataset
In this section, we compare commonly used image-caption datasets with the SynthVLM synthetic dataset. The synthetic data offers high image quality, excellent image-text alignment, superior machine ratings, and robust data privacy protection.
High Image Quality. As illustrated in Figure 1, SynthVLM significantly advances image quality by providing images at a resolution of 1024x1024 pixels. This high resolution addresses the common issue of inadequate image quality in existing datasets, thereby supplying high-quality image-caption pairs invaluable for VLLMs. Furthermore, Curated-SynthVLM effectively mitigates issues such as watermarks and advertisements. For image-caption alignment, Curated-SynthVLM leverages the SoTA SDXL to generate images, ensuring that generated images closely correspond to the provided captions.
Excellent Image-Text Alignment. As shown in Table 2, the generated SynthVLM dataset exhibits a higher CLIPScore. By selecting curated image-text pairs of higher quality, Curated-SynthVLM achieves an even higher CLIPScore, surpassing COCO-Caption, BLIP-LCS, and ShareGPT4V. This demonstrates the excellent alignment of our data.
Table 3. We employed GPT4-Vision and InternVL to vote on the match between each caption and its corresponding generated image and raw image. The results demonstrate that the generated images align more closely with the captions.
Sample Model Image-gen win Image-raw win
1K GPT4-Vision 633 367
1K InternVL 692 308
Excellent Machine Rating. Since our data will be used for VLLMs training, we use VLLMs to evaluate the data quality. We selected 1K image-caption pairs and submitted the caption along with the synthetic image and the original image. We used GPT-4 Vision and Intern-VL as the judge model and requested it to select the pair that exhibited higher alignment. The specific prompt used for this evaluation is illustrated in Figure 4. The results, presented in Table 3, demonstrate that images generated have better alignment with the caption.
Protect Data Privacy. A notable advantage of our dataset is its exclusive reliance on captions, which mitigates data privacy concerns. By not using existing images, we ensure that no sensitive or private information associated with those images is compromised. This approach adheres to data privacy best practices, ensuring that our dataset maintains both high quality and ethical integrity.
3.4. SynthVLM
In this section, we utilize the synthetic dataset in section 3.3 to pre-train the Vision Language model. We adopt the same widely-used model architecture summarized in (Zhang et al., 2024; Liu et al., 2023b), as depicted in Figure 5.
Pre-training Stage
As illustrated in Figure 5(a), we train the projector during the pre-training stage to achieve alignment between the image and text modalities. The SynthVLM dataset described in Section 3.3 is utilized for this purpose.
SFT Stage
As shown in Figure 5(b), we further train the projector along with the LLM during the SFT stage to enhance visual understanding capabilities. For this stage, we utilize the commonly used LLaVA 665k dataset from (Liu et al., 2023a).
Through these two training stages, we successfully developed the SynthVLM model. SynthVLM is efficient, utilizing only 100k pre-training data while also protecting privacy by leveraging synthetic data. Additionally, SynthVLM provides a new paradigm for effective alignment between modalities in Vision Language Models using synthetic data.
Figure 5. We introduce the pre-training model structure in Figure (a). Using synthetic data, we pre-train the projector to align the image and text modalities. As shown in Figure (b), we subsequently use LLaVA 665k data to fine-tune the projector and the LLM.
4. Experiment
In this section, we first introduce the experimental setups. We then aim to answer the following questions to verify the effectiveness, efficiency, and privacy protection of our proposed SynthVLM: Q1: Can our SynthVLM achieve SoTA performance compared to previous SoTA methods? Q2: Can our SynthVLM have better image-text alignment compared to previous methods? Q3: How efficient is our SynthVLM compared to previous methods? Q4: Can SynthVLM protect privacy while achieving SoTA performance? Q5: Do we need the generate module and the data quality selection module to enhance model performance?
4.1. Experimental Settings
Datasets.
We utilized the 558k pre-training dataset and the 665k SFT dataset from LLaVA, in addition to the synthetic 100k dataset in section 3.3 for training our SynthVLM.
Models.
For the image generation model, we select SDXL, and SDXL’s steps are set to 60. We employed the well-known LLaVA 1.5 model configured with 13 billion parameters for its robust visual understanding capabilities. For the encoder, we chose the CLIP 336 with a 14-patch configuration, and for the language model, the Vicuna v1.5 configured with 13 billion parameters was selected.
Baselines.
The LLaVA 1.5111https://github.com/haotian-liu/LLaVA model was reproduced using configurations sourced from the official repository to establish a baseline. We opted for the LLaVA model trained with Vicuna v1.5 7B and 13B models as the baseline (Liu et al., 2023b).
Benchmarks.
We select benchmarks for both visual understanding and language understanding. For visual understanding, we choose SQA_Img, MMVet, VizWiz, VQAv2, GQA, MME, and PoPE for a comprehensive evaluation. For pure text benchmarks, we select MMLU and SQA to assess language understanding abilities.
Table 4. Comparison of SynthVLM and LLaVA using the same model structure. We can see SynthVLM outperforms LLaVA on all the evaluation benchmarks.
Models LLM SQA SQA_Img MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
Baseline Vicuna-1.5-7B 69.3 67.3 30.5 49.9 78.7 62.5 65.3 1484.8 315.6 86.0 36.3
Synth Select 100k Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
Baseline Vicuna-1.5-13B 74.2 71.0 35.0 53.6 80.0 63.0 67.7 1531.3 294.5 86.9 52.4
Synth Select 100k Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
Table 5. Result comparison of MMLU shows that with the synthetic 100k data, our SynthVLM outperforms LLaVA in pure language tasks. This demonstrates the effectiveness of the synthetic data in modality alignment.
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
Baseline Vicuna-1.5-7B 69.3 36.3 28.6 33.4 39.5 44.5
Synth Select 100k Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
Baseline Vicuna-1.5-13B 74.2 52.4 41.9 45.8 62.9 61.8
Synth Select 100k Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
4.2. Synthetic Data Achieves SoTA Performance
To address Q1, we selected LLaVA 1.5 with the CLIP 336 encoder and Vicuna v1.5 7B and 13B as the base model. We used the synthetic 100k dataset to pretrain the LLaVA 1.5 model and subsequently fine-tuned it with LLaVA 665k SFT data. The trained model is denoted as ”Synth Select 100k.” We then compared it with the LLaVA 1.5 model pre-trained on 558k data and fine-tuned on the same 665k SFT data from the official repository, referred to as ”Baseline”
From Table 4, it is evident that Synth Select 100k outperforms the Baseline across all evaluation benchmarks on both the 7B and 13B model. Specifically, Synth Select 100k achieves SoTA results on visual benchmarks such as SQA_Img, MMVet, VizWiz, VQAv2, GQA, MME, MMB, and PoPE. Furthermore, SynthVLM also excels in pure language benchmarks, demonstrating superior performance in SQA and MMLU, thus showcasing its comprehensive capabilities in both vision and language tasks.
4.3. Effective Vision Language Alignment
To address Q2, we utilize pure language ability to demonstrate the model’s alignment. During the SFT stage, the LLM can be adjusted to align with images; hence, better alignment can preserve the LLM’s pure language ability (Wang et al., 2023). We select the MMLU and SQA benchmarks for their comprehensive language understanding capabilities.
From Table 5, it is evident that our Synth Select 100k model outperforms the Baseline on all MMLU benchmarks and SQA benchmarks.
These results demonstrate the strong alignment capability of our synthetic data. Additionally, this provides a new paradigm for effective visual understanding model modality alignment using generated data. During pre-training, it is common to train on all available data due to uncertainty about data selection. Here, we offer 100k high-quality synthetic data as a benchmark for selecting aligned generated data efficiently.
Table 6. Comparison of data utilization for generating image-caption pairs. This indicates that our SynthVLM have superior efficiency compared to other methods.
Methods SynthVLM LLaVA w/o selection
Dataset Number (k) 100 558 1000
Data Usage 33MB 27GB 330MB
4.4. Efficient Vision Language Alignment
To address Q3, we examine the computational resource usage during training and evaluate data utilization efficiency for generating image-caption pairs.
Computational Resource Usage
As shown in Table 6, by integrating a data selection module, our approach utilizes only 19% of the LLAVA data and 10% of the original synthetic data while achieving SoTA performance. This demonstrates that our data selection method can reduce computational usage by more than 80%.
Data Utilization Efficiency
Table 6 compares the data utilization of our proposed method with the conventional BLIP-LCS method, which employs the BLIP model for caption generation. Our approach is significantly more efficient, requiring only 330 MB for captions to generate 1,000,000 image-caption pairs, compared to the traditional methods, which may exceed 50 GB for images. This substantial difference highlights the efficiency of relying solely on captions rather than images in generating image-caption pairs.
Overall, our method efficiently aligns image and text modalities, demonstrating strong potential for effective modality alignment. Additionally, acquiring captions is considerably less resource-intensive than obtaining images. By integrating existing large language models (LLMs), it is feasible to fully automate the generation of both captions and images, further enhancing the efficiency of dataset construction in VLLMs.
4.5. Privacy Protection Pre-training
To address Q4, we compare the synthetic image and the original image in Figure 6 and Figure 7. We can see synthetic data offers significant advantages in protecting data privacy.
As illustrated in Figure 6, synthetic image (a) effectively avoids privacy issues by not representing real human faces, while original image (b) contains human faces, potentially leading to privacy concerns. Similarly, in Figure 7, synthetic images in (a) show vehicles and tickets without revealing real license plates and ticket information, ensuring privacy protection. In contrast, original images in (b) display actual license plates and ticket information, which can potentially lead to privacy issues.
By utilizing generative models like DDPM, synthetic data can be created with similar statistical properties to real data without involving actual personal information. This provides a secure pathway for data sharing, model training, and analysis, helping to comply with privacy regulations, protecting user privacy, and simultaneously advancing the fields of artificial intelligence and machine learning.
Table 7. Ablation study of visual understanding ability and pure language ability. The results demonstrate that removing either the data generation or data selection module results in a performance drop.
Models LLM SQA SQA_Img MMVet VizWiz VQAv2 GQA MMB MMEP MMEC PoPE MMLU
Synth Select 100k Vicuna-1.5-7B 70.4 68.9 32.2 49.3 79.4 63.1 66.8 1518.5 345.7 87.0 41.2
w/o generation 100k Vicuna-1.5-7B 69.3$\downarrow$ 67.0$\downarrow$ 31.2$\downarrow$ 46.8$\downarrow$ 79.3$\downarrow$ 62.9$\downarrow$ 66.2$\downarrow$ 1488.8$\downarrow$ 327.5$\downarrow$ 86.2$\downarrow$ 39.1$\downarrow$
w/o selection 100k Vicuna-1.5-7B 69.9$\downarrow$ 67.7$\downarrow$ 30.2$\downarrow$ 50.2 79.1$\downarrow$ 62.2$\downarrow$ 63.5$\downarrow$ 1421.7$\downarrow$ 301.8$\downarrow$ 87.3 40.6$\downarrow$
Synth Select 100k Vicuna-1.5-13B 74.9 72.5 35.0 55.9 80.0 63.5 68.3 1573.0 316.1 88.4 54.6
w/o generation 100k Vicuna-1.5-13B 73.6$\downarrow$ 71.4$\downarrow$ 33.0$\downarrow$ 53.6$\downarrow$ 80.0 63.4$\downarrow$ 67.5$\downarrow$ 1514.3$\downarrow$ 295.7$\downarrow$ 88.2$\downarrow$ 53.6$\downarrow$
w/o selection 100k Vicuna-1.5-13B 74.1$\downarrow$ 70.5$\downarrow$ 35.6 53.2$\downarrow$ 79.7$\downarrow$ 63.1$\downarrow$ 67.5$\downarrow$ 1512.7$\downarrow$ 303.2$\downarrow$ 86.9$\downarrow$ 53.0$\downarrow$
Table 8. Ablation study of modality alignment. The results demonstrate that removing either the data generation or data selection module results in a performance drop.
Models LLM SQA MMLU
Avg STEM Humanities Social Sciences Other
Synth Select 100k Vicuna-1.5-7B 70.4 41.2 31.7 37.4 47.0 50.2
w/o generation Vicuna-1.5-7B 69.3$\downarrow$ 39.1$\downarrow$ 30.0$\downarrow$ 36.6$\downarrow$ 43.1$\downarrow$ 47.3$\downarrow$
w/o selection Vicuna-1.5-7B 69.9$\downarrow$ 40.6$\downarrow$ 30.8$\downarrow$ 37.2$\downarrow$ 45.3$\downarrow$ 48.9$\downarrow$
Synth Select 100k Vicuna-1.5-13B 74.9 54.6 45.0 49.3 64.0 62.2
w/o generation Vicuna-1.5-13B 74.1$\downarrow$ 53.6$\downarrow$ 43.5$\downarrow$ 48.2$\downarrow$ 63.1$\downarrow$ 61.8$\downarrow$
w/o selection Vicuna-1.5-13B 73.6$\downarrow$ 53.0$\downarrow$ 42.9$\downarrow$ 46.8$\downarrow$ 63.8$\downarrow$ 61.3$\downarrow$
Figure 6. From (a), it is evident that synthetic images can avoid representing real human faces, while (b) contains human faces, potentially leading to privacy issues.
Figure 7. From (a), it is evident that synthetic images can avoid displaying real license plates and ticket information. In contrast, (b) contains actual license plates and ticket information, which can potentially lead to privacy issues.
4.6. Ablation Study
To address Q5, we conducted the following ablation study. Specifically, we conducted an ablation study where we removed the data generation module and the data selection module separately to evaluate their individual contributions to the effectiveness of our data generation pipeline.
Excluding Data Generation Module
The exclusion of the data generation module significantly impacts the model’s performance, as illustrated in Tables 7 and 8, labeled as ”w/o generation 100k”. The variant without this module demonstrates markedly lower accuracy across all benchmarks. These results emphasize the crucial role of the data generation process in sustaining the high performance of the SynthVLM model. This also underscores SynthVLM’s potential in constructing highly aligned datasets.
Excluding Data Selection Module
The absence of the data selection module similarly leads to a noticeable decline in performance, indicated as ”w/o selection 100k” in Tables 7 and 8. Given the inherent randomness of diffusion models, which inevitably generate some low-quality images, the data selection module is crucial for removing these subpar elements.
Overall, the ablation study highlights the critical role of data generation and data selection modules in SynthVLM. These experiments provide valuable insights into the contributions of each module, guiding future improvements and optimizations of the SynthVLM model.
5. Conclusion
In recent years, with the development of VLLMs, data generation has become increasingly important. Synthetic images are crucial for LLMs due to the lack of high-quality data for extensive training. In this paper, we propose a new image generation pipeline for generating VLLMs pre-training data, providing a new paradigm for image generation for VLLMs. Remarkably, the synthetic data trained SynthVLM model outperforms the baseline using only 18% synthetic data. Additionally, it achieves SoTA alignment ability efficiently. Furthermore, SynthVLM protects data privacy by using synthetic datasets.